# AdaRank: Adaptive Rank Pruning for Enhanced Model Merging

### v2

@@ 2 Preliminaries
where $\text{SVD}_{k}$ denotes the Singular Value Decomposition with low-rank approximation on top-$k$ singular components. Unlike element-wise modification, truncating the singular components of the task vector retains its core low-dimensional structure of the original matrix, thereby preserving the rich information of individually fine-tuned task vectors. Meanwhile, maintaining the rank of task vectors $k$ small helps to reduce the potential of inter-task interference across tasks (Lee et al., 2025).
@@ 3 Analysis on SVD of Task Vectors
Building on the low-rank paradigm, we seek an alternative direction for enhancing model merging. Specifically, we scrutinize the common practice of truncating each task vector to its top-k{k} singular components, posing two central questions: (1) Is choosing the top singular components always beneficial for model merging? (2) Is enforcing a fixed rank across tasks and layers desirable for model merging? In this section, we conduct an empirical analysis to investigate these questions, revealing the limitations of naive top-$k$ truncation.
@@
While it is guaranteed that the top-$k$ singular components yield the best low-rank approximation in terms of reconstruction error for a single matrix  (Eckart & Young, 1936), this optimality does not necessarily transfer to multi-task model merging. In a single-task scenario, these top components—characterized by large singular values—effectively minimize the loss for that particular task. However, in a multi-task setting, these top components can also exhibit significant interference across tasks, potentially degrading the merging performance.
@@
For a deeper understanding of this phenomenon, we analyze the loss changes of individual tasks exemplified by the case where the excluded task $i$ is MNIST (LeCun, 1998) (Figure 1 (c)). We observe that adding a singular component from the MNIST benefits semantically aligned tasks (e.g., SVHN (Netzer et al., 2011), digit classification task), whereas dissimilar tasks (e.g., DTD (Cimpoi et al., 2014), a texture classification task) experience increases in a net loss. These varying interactions lead to a complex interplay of losses for each singular component. In particular, top components with large singular values $\sigma_{ir}$ can have a more pronounced negative impact on dissimilar tasks, increasing their net loss. Overall, these results suggest that top singular components are not always optimal for multi-task model merging.
@@
Another significant limitation lies in the use of a fixed top-$k$ truncation across diverse tasks and model layers. In Figure 2, we quantify the intrinsic rank of the task vectors by measuring the number of singular components required to preserve 95% of the total variance, or spectral energy (i.e., the sum of squared singular values) (Jolliffe, 2011; Raschka & Mirjalili, 2019), which captures the effective linear dimensionality of the matrix. The intrinsic rank varies considerably across task vectors from different models. Consistent with the finding that the intrinsic dimension of neural network representations increases with task complexity (Li et al., 2018), our experiment supports that this trend is also observed in task vectors, revealing a strong correlation between intrinsic rank and task demands. For instance, task vectors from models fine-tuned with SUN397 (Xiao et al., 2016) (397 classes) require a higher rank compared to those from simpler tasks like MNIST (LeCun, 1998) or SVHN (Netzer et al., 2011). Moreover, task vectors exhibit pronounced rank variation across layers (see Figure 2). Early layers, which capture task-agnostic features (Krizhevsky et al., 2012; Raghu et al., 2017; Papyan et al., 2020), show higher ranks with lower variance, reflecting shared information across tasks. In contrast, later layers, encoding task-specific representations, demonstrate greater rank variability and lower overall ranks. This divergence of rank among tasks and layers highlights the challenge of using a fixed top-$k$ truncation for model merging, since we may either discard important components that are critical for some tasks or keep unnecessary components that cause interference between tasks.
@@ 6.1 Experimental Setup
To demonstrate the generality of our method, we integrate AdaRank into two different baselines: Task Arithmetic (Ilharco et al., 2023) with SVD and CART (Choi et al., 2024), which correspond to Equation 2 with the difference in setting $\theta_{0}$ as a pre-trained model or weight averaging, respectively. For optimization, we initialize the binary matrix $B$ such that Equation 5 is initialized to Equation 1 for Task Arithmetic and to Equation 2 for CART. Next, we employ an additional best-performing SVD-based merging method, TSV-M (Gargiulo et al., 2025). For TSV-M, a whitening transform is applied to $U_{i}$ and $V_{i}$ matrices from Equation 5 (see Appendix A for details). Since the layer-wise coefficient $\lambda^{l}$ can be learned jointly with the binary mask $B$ via TTA as in AdaMerging (Yang et al., 2023), we learn both of them jointly in our default setting, with the effect from each $\lambda$ and $B$ is analyzed in Section 6.3. The specific initial values for $\lambda$ and $B$ are set according to the best-performing configurations described in their respective papers, with further details provided in Appendix B.1.
@@ 6.2 Main Results
Following standard experiment protocol (Wang et al., 2024), we evaluate our method by merging two Vision Transformer (Dosovitskiy et al., 2020) backbones of CLIP (Radford et al., 2021): ViT-B/32 and ViT-L/14, which are fine-tuned on 8, 14, and 20 image classification tasks. The selected tasks cover a diverse range of domains and class counts, from fine-grained image classification on Flowers102 (Nilsback & Zisserman, 2008) to character recognition on KMNIST (Clanuwat et al., 2018). We provide detailed dataset lists and descriptions in Appendix B.2.
@@
We provide a detailed analysis of the number of learnable parameters and the training time of AdaRank in Appendix C.2. Compared to AdaMerging (Yang et al., 2023), AdaRank incurs a marginal computational overhead while achieving significant improvements.
@@ 6.3 Analysis and Ablation
Effect of Selecting Bottom Singular Components.
In Figure 5 (a), we show the cumulated count of singular component indices selected by AdaRank, compared to fixed top-$k$. It indicates that AdaRank not only prunes the top singular components but also frequently selects the bottom components. To quantify the performance gain from selecting the bottom components, we conduct an ablation study on the selection range. Starting from the best-performing top-$16\%$ truncation, we compare AdaRank applied only on top-$k$ range (i.e, $B_{ir}$ = 0 always for $r>k$) with that used on the full range. As shown in Table 4, this variant also presents a significant performance gain, showing that removing conflicting top components is critical for mitigating task interference, as discussed in Section 3. However, AdaRank with a full range gains further by incorporating components from the indices outside the 16% range. We conjecture that bottom components provide valuable gains for their own task while introducing less inter-task interference than top components, aligning with the observations that they often correlate with fine details (Dabov et al., 2007; Candes & Plan, 2010; Lee et al., 2025).
Adaptive Rank Assignment Across Tasks and Layers.
@@
Ablation on Learnable Components.
@@ Appendix A Algorithm Details
We summarize the procedure of AdaRank in Algorithm 1. Then, we explain how to initialize the layer-wise task vectors and apply SVD for different baseline methods in Algorithm 2.
@@
1:Pretrained Weight $\theta_{0}$, Fine-Tuned Weights $\{\theta_{i}\}_{i=1}^{T}$, Test Dataset $\{\mathcal{D}_{i}\}$, Task Coefficient $\{\lambda^{l}\}$, Adaptation Learning Rate $\eta$, Method
2:
[…]
4:Initialize mask parameters $B=\{B_{i}^{l}\}_{i=1,l=1}^{T,L}$.
5:
6:while not converged do
7:  Construct merged model $\theta(B)$ using $\{U_{i}^{l},\Sigma_{i}^{l},V_{i}^{l}\}$ and $B$ via Equation 5.
8:  Compute total entropy loss $\mathcal{L}\leftarrow\sum_{i=1}^{T}\mathbb{E}_{x_{i}\sim\mathcal{D}_{i}}[H(f(\theta(B),x_{i}))]$.
9:  Update parameters $B\leftarrow B-\eta\nabla_{B}\mathcal{L}$ using Straight-Through Estimator.
10:end while
11:
12:Construct final merged model $\theta^{*}$ using finalized mask $B^{*}$ via Equation 5.
13:return $\theta^{*}$.
Algorithm 2 SVD Components Preparation
1:procedure PrepareSVDComponents($\{\theta_{i}\}_{i=1}^{T},\theta_{0},\text{Method}$)
2:
3:  for all task $i=1,\dots,T$ do
4:   if Method is Task Arithmetic or TSV-M then
5:     Calculate task vector: $\tau_{i}\leftarrow\theta_{i}-\theta_{0}$
6:   else if Method is CART then
7:     Calculate average of fine-tuned weights: $\bar{\theta}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\theta_{j}$
8:     Calculate task vector: $\tau_{i}\leftarrow\theta_{i}-\bar{\theta}$
9:   end if
10:  end for
11:
12:  for all layer $l=1,\dots,L$ do
13:   if Method is TSV-M then
14:     for all task $i=1,\dots,T$ do
15:      $U_{i}^{l},\Sigma_{i}^{l},V_{i}^{l}\leftarrow\text{SVD}(\tau_{i}^{l})$
16:     end for
17:     Concatenate into single matrix: $U^{l}\leftarrow[U_{1}^{l}|\dots|U_{T}^{l}]$,  $V^{l}\leftarrow[V_{1}^{l,}|\dots|V_{T}^{l}]$
18:     Apply whitening transformation: $U_{\perp}^{l}\leftarrow\text{whitening}(U^{l})$,  $V_{\perp}^{l}\leftarrow\text{whitening}(V^{l})$
19:     Split $U_{\perp}^{l}$ and $V_{\perp}^{l}$ back into task-specific matrices $\{U_{i}^{l}\}_{i=1}^{T}$ and $\{V_{i}^{l}\}_{i=1}^{T}$.
20:   else
21:     for all task $i=1,\dots,T$ do
[…]
23:      $U_{i}^{l},\Sigma_{i}^{l},V_{i}^{l}\leftarrow\text{SVD}(\tau_{i}^{l})$
24:     end for
25:   end if
26:  end for
27:  return $\{U_{i}^{l},\Sigma_{i}^{l},V_{i}^{l}\}_{i=1,l=1}^{T,L}$
28:end procedure
@@ B.1 Implementation Details
For test-time adaptation (TTA), we share the same hyperparameters and TTA budget for all the adaptive merging and router-based merging baselines (Yang et al., 2023; Wei et al., 2025b; Lu et al., 2024; Tang et al., 2024b). We followed the settings from AdaMerging (Yang et al., 2023), using the Adam (Kingma & Ba, 2015) optimizer with a learning rate of $1\times 10^{-3}$ for vision model merging and $5\times 10^{-5}$ for language model merging. Both configurations used Adam betas of $(0.9,0.999)$ and a test batch size of $16$ per task.
@@ B.2 Dataset Description
  • Flowers102 (Nilsback & Zisserman, 2008) is a fine-grained flower classification dataset with 102 categories, where each class contains between 40 and 258 images.
@@ B.3 Datasets for Language Tasks
  • MRPC (Dolan & Brockett, 2005) is a binary classification task to determine if two sentences are semantically equivalent. The standard GLUE split consists of 3.7k training and 1.7k test pairs.

### v3

# arXiv:2503.22178v3[cs.LG] 28 Feb 2026

@@ 2 Preliminaries
where $\text{SVD}_{k}$ denotes the Singular Value Decomposition with low-rank approximation on top-$k$ singular components. Unlike element-wise modification, truncating the singular components of the task vector retains the core low-dimensional structure of the original matrix, thereby preserving the rich information of individually fine-tuned task vectors. Meanwhile, maintaining the rank of task vectors $k$ small helps to reduce the potential of inter-task interference across tasks (Lee et al., 2025).
@@ 3 Analysis on SVD of Task Vectors
Building on the low-rank paradigm, we seek an alternative direction for enhancing model merging. Specifically, we scrutinize the common practice of truncating each task vector to its top-k{k} singular components, and pose two central questions: (1) Is choosing the top singular components always beneficial for model merging? (2) Is enforcing a fixed rank across tasks and layers desirable for model merging? In this section, we conduct an empirical analysis to investigate these questions, revealing the limitations of naive top-$k$ truncation.
@@
While it is guaranteed that the top-$k$ singular components yield the best low-rank approximation in terms of reconstruction error for a single matrix (Eckart and Young, 1936), this optimality does not necessarily transfer to multi-task model merging. In a single-task scenario, these top components—characterized by large singular values—effectively minimize the loss for that particular task. However, in a multi-task setting, these top components can also exhibit significant interference across tasks, potentially degrading the merging performance.
@@
For a deeper understanding of this phenomenon, we analyze the loss changes of individual tasks exemplified by the case where the excluded task $i$ is MNIST (LeCun, 1998) (Figure 1 (c)). We observe that adding a singular component from the MNIST benefits semantically aligned tasks (e.g., SVHN (Netzer et al., 2011), digit classification task), whereas dissimilar tasks (e.g., DTD (Cimpoi et al., 2014), a texture classification task) experience increases in net loss. These varying interactions lead to a complex interplay of losses for each singular component. In particular, top components with large singular values $\sigma_{ir}$ can have a more pronounced negative impact on dissimilar tasks, increasing their net loss. Overall, these results suggest that top singular components are not always optimal for multi-task model merging. For additional extensive analysis on the singular components of task vectors, see Appendix C.1.
@@
Another significant limitation lies in the use of a fixed top-$k$ truncation across diverse tasks and model layers. In Figure 2, we quantify the intrinsic rank of the task vectors by measuring the number of singular components required to preserve 95% of the total variance, or spectral energy (i.e., the sum of squared singular values) (Jolliffe, 2011; Raschka and Mirjalili, 2019), which captures the effective linear dimensionality of the matrix. The intrinsic rank varies considerably across task vectors from different models. Consistent with the finding that the intrinsic dimension of neural network representations increases with task complexity (Li et al., 2018), our experiment supports that this trend is also observed in task vectors, revealing a strong correlation between intrinsic rank and task demands. For instance, task vectors from models fine-tuned with SUN397 (Xiao et al., 2016) (397 classes) require a higher rank compared to those from simpler tasks like MNIST (LeCun, 1998) or SVHN (Netzer et al., 2011). Moreover, task vectors exhibit pronounced rank variation across layers (see Figure 2). Early layers, which capture task-agnostic features (Krizhevsky et al., 2012; Raghu et al., 2017; Papyan et al., 2020), show higher ranks with lower variance, reflecting shared information across tasks. In contrast, later layers, encoding task-specific representations, demonstrate greater rank variability and lower overall ranks. This divergence of rank among tasks and layers highlights the challenge of using a fixed top-$k$ truncation for model merging, since we may either discard important components that are critical for some tasks or keep unnecessary components that cause interference between tasks.
@@ 6.1 Experimental Setup
To demonstrate the generality of our method, we integrate AdaRank into two different baselines: Task Arithmetic (Ilharco et al., 2023) with SVD and CART (Choi et al., 2024), which correspond to Equation 2 with the difference in setting $\theta_{0}$ as a pre-trained model or weight averaging, respectively. For optimization, we initialize the binary matrix $B$ such that Equation 5 is initialized to Equation 1 for Task Arithmetic and to Equation 2 for CART. Additionally, we employ TSV-M (Gargiulo et al., 2025) and Iso-CTS (Marczak et al., 2025), another top-performing SVD-based merging method. For TSV-M, a whitening transform is applied to $U_{i}$ and $V_{i}$ matrices from Equation 5 (see Appendix A for details). Since the layer-wise coefficient $\lambda^{l}$ can be learned jointly with the binary mask $B$ via TTA as in AdaMerging (Yang et al., 2023), we learn both of them jointly in our default setting, with the effect from each $\lambda$ and $B$ is analyzed in Section 6.3. The specific initial values for $\lambda$ and $B$ are set according to the best-performing configurations described in their respective papers, with further details provided in Appendix B.1.
@@ 6.2 Main Results
Following standard experiment protocol (Wang et al., 2024), we evaluate our method by merging two Vision Transformer (Dosovitskiy et al., 2020) backbones of CLIP (Radford et al., 2021): ViT-B/32 and ViT-L/14, which are fine-tuned on 8, 14, and 20 image classification tasks. The selected tasks cover a diverse range of domains and class counts, from fine-grained image classification on Flowers102 (Nilsback and Zisserman, 2008) to character recognition on KMNIST (Clanuwat et al., 2018). We provide detailed dataset lists and descriptions in Appendix B.2.
@@
Method Total Params. Learnable Params. SVD Time TTA Time Performance
AdaMerging (ViT-B/32) 907.6M 2,440 - $\sim$10.3 min 80.1%
AdaMerging (ViT-L/14) 2.46B 3,592 - $\sim$36.7 min 90.8%
@@ 6.3 Analysis and Ablation
Effect of Selecting Bottom Singular Components.
In Figure 4 (a), we show the cumulated count of singular component indices selected by AdaRank, compared to fixed top-$k$. It indicates that AdaRank not only prunes the top singular components but also frequently selects the bottom components. To quantify the performance gain from selecting the bottom components, we conduct an ablation study on the selection range. Starting from the best-performing top-$16\%$ truncation, we compare AdaRank applied only on top-$k$ range (i.e, $B_{ir}$ = 0 always for $r>k$) with that used on the full range. As shown in Table 6, this variant also presents a significant performance gain, showing that removing conflicting top components is critical for mitigating task interference, as discussed in Section 3. However, AdaRank with a full range gains further by incorporating components from the indices outside the 16% range. We conjecture that bottom components provide valuable gains for their own task while introducing less inter-task interference than top components, aligning with the observations that they often correlate with fine details (Dabov et al., 2007; Candes and Plan, 2010; Lee et al., 2025).
@@
Adaptive Rank Assignment Across Tasks and Layers.
Ablation on Learnable Components.
Robustness to the Amount of Test Data.
Lastly, we examine the robustness of AdaRank with respect to the amount of test data available for TTA. Table 7 reports the performance when varying the proportion of accessible test data from $1\%$ to $100\%$. Even with only $1\%$ of the test set, AdaRank achieves gains of $12.0\%$ and $3.3\%$ on the 8- and 20-task benchmarks, respectively, surpassing AdaMerging adapted with the full test set. This trend holds consistently across both TA and CART baselines, demonstrating that AdaRank is robust to limited data availability and it offers an effective solution for real-world scenarios. For results with NLP tasks, see Table 10 in appendix.
@@
Table 7: Performance of AdaRank with varying amounts of test data on ViT-B/32. The leftmost columns report the baseline without TTA (0%) and AdaMerging using the full test set for reference.
Method No TTA AdaMerging AdaRank
Amount of Test Data 0% 100% 1% 5% 10% 50% 100%
8 tasks TA 69.2 80.1 81.2 84.2 85.2 87.2 87.9
CART 65.4 76.7 76.8 78.2 79.0 81.3 82.1
14 tasks TA 61.0 69.2 70.0 71.3 73.0 76.7 81.4
CART 84.7 85.9 86.1 87.1 87.7 88.6 89.2
20 tasks TA 79.5 82.3 82.8 84.4 85.3 85.8 86.2
CART 76.8 82.7 82.7 84.3 85.2 85.8 86.4
@@ 7 Conclusion
Acknowledgements
This work is supported by the National Research Foundation of Korea (RS-2024-00351212 and RS-2024-00436165), the Institute of Information & communications Tech- nology Planning & Evaluation (IITP) (RS-2024-00509279, RS-2022-II220926 , RS-2022-II220959, and RS-2019-II190075), funded by Korea Government (MSIT).
@@ Appendix A Algorithm Details
We summarize the full procedure of AdaRank in Algorithm 1, including the SVD preparation step for each baseline method and the test-time adaptation loop.
@@
// SVD Components Preparation
Require: Pretrained weight $\theta_{0}$, fine-tuned weights $\{\theta_{i}\}_{i=1}^{T}$, Method
for task $i=1$ to $T$ do
  if Method is TA or TSV-M then
   $\tau_{i}\leftarrow\theta_{i}-\theta_{0}$
  else if Method is CART then
   $\bar{\theta}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\theta_{j}$
   $\tau_{i}\leftarrow\theta_{i}-\bar{\theta}$
  end if
end for
for layer $l=1$ to $L$ do
  for task $i=1$ to $T$ do
   $U_{i}^{l},\Sigma_{i}^{l},V_{i}^{l}\leftarrow\text{SVD}(\tau_{i}^{l})$
  end for
  if Method is TSV-M then
   $U^{l}\leftarrow[U_{1}^{l}|\dots|U_{T}^{l}]$,  $V^{l}\leftarrow[V_{1}^{l}|\dots|V_{T}^{l}]$
   $U_{\perp}^{l}\leftarrow\text{whitening}(U^{l})$,  $V_{\perp}^{l}\leftarrow\text{whitening}(V^{l})$
   Split $U_{\perp}^{l},V_{\perp}^{l}$ back into per-task matrices $\{U_{i}^{l}\}_{i=1}^{T},\{V_{i}^{l}\}_{i=1}^{T}$
  end if
end for
// Test-Time Adaptation
Require: Test datasets $\{\mathcal{D}_{i}\}_{i=1}^{T}$, task coefficients $\{\lambda^{l}\}$, learning rate $\eta$
Initialize: Mask parameters $B=\{B_{i}^{l}\}_{i=1,l=1}^{T,L}$
repeat
  Construct merged model $\theta(B)$ via Equation 5
  $\mathcal{L}\leftarrow\sum_{i=1}^{T}\mathbb{E}_{x_{i}\sim\mathcal{D}_{i}}[H(f(\theta(B),x_{i}))]$
  Update $B\leftarrow B-\eta\nabla_{B}\mathcal{L}$ using Straight-Through Estimator (Appendix A)
until converged
 Construct final merged model $\theta^{*}$ using finalized $B^{*}$ via Equation 5
Return $\theta^{*}$
@@ B.1 Implementation Details
For test-time adaptation (TTA), we share the same hyperparameters and TTA budget for all the adaptive merging and router-based merging baselines (Yang et al., 2023; Wei et al., 2025b; Lu et al., 2024; Tang et al., 2024b). We followed the settings from AdaMerging (Yang et al., 2023), using the Adam (Kingma and Ba, 2015) optimizer with a learning rate of $1\times 10^{-3}$ for vision model merging and $5\times 10^{-5}$ for language model merging. Both configurations used Adam betas of $(0.9,0.999)$ and a test batch size of $16$ per task.
@@ B.2 Dataset Description
  • Flowers102 (Nilsback and Zisserman, 2008) is a fine-grained flower classification dataset with 102 categories, where each class contains between 40 and 258 images.
@@ B.3 Datasets for Language Tasks
  • MRPC (Dolan and Brockett, 2005) is a binary classification task to determine if two sentences are semantically equivalent. The standard GLUE split consists of 3.7k training and 1.7k test pairs.
@@
Figure 6: Net change in single-task and multi-task losses for (a) RoBERTa (Liu et al., 2019) and (b) GPT-2 (Radford et al., 2019), respectively, when each singular component of a target task vector is individually added to a model merged with full-rank vectors from other tasks. Task vectors are defined as deviation from pre-trained weight as Task Arithmetic Ilharco et al. (2023). Only Top-10% indices are presented for clarity.
@@ Appendix C Additional Results and Analyses
C.1 Extended Analysis on SVD of Task Vectors
Loss analysis on different models.
First, we verify that the trend in loss change induced by singular component addition is consistent across different scenarios. In Figure 6, we replicate the loss analysis on RoBERTa (Liu et al., 2019) and GPT-2 (Radford et al., 2019), where we individually add a singular component of a single task vector to a model merged with full-rank task vectors from the other tasks. Regardless of the change in architecture, we observe a similar trend as in the ViT: top components strongly benefit their own task loss, but some of those components also increase the net multi-task loss. This confirms that our finding is a general phenomenon across diverse model merging scenarios, not limited to ViT with image classification tasks.
Condition for interference.
To understand why such components appear, we first analyze the change in multi-task loss $L(\theta)=\sum_{t=1}^{T}L_{t}(\theta)$ induced by adding a singular component. Consider adding an arbitrary singular component from task $i$, denoted as $\mathbf{s}_{i}=\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$, to the model parameter $\theta$. We rewrite the loss change $\Delta L=L(\theta+\mathbf{s}_{i})-L(\theta)$ in terms of the direction of the component $\bar{\mathbf{s}}_{i}=\mathbf{u}_{i}\mathbf{v}^{\top}_{i}$ and the singular value $\sigma_{i}$, using a second-order Taylor expansion:
$\Delta L\approx\sigma_{ir}\nabla L^{\top}\mathbf{\bar{s}}_{i}+\frac{1}{2}\sigma_{i}^{2}\mathbf{\bar{s}}_{i}^{\top}\mathbf{H}\mathbf{\bar{s}}_{i},$ (7)
where $\mathbf{H}$ is the Hessian of the loss at $\theta$. The value of $\Delta L$ is determined by two factors:
  1. 1.
    Directional Alignment: The alignment between the singular component direction $\mathbf{\bar{s}}_{ir}$ and the multi-task loss landscape (gradient and Hessian)
  2. 2.
    Spectral Scale: The magnitude of the singular value $\sigma_{i}$, which scales each first and second-order term.
We showed the value of $\Delta L$ as “Multi-Task Loss Change” plot in Figure 1, where it implicitly contains both the effects from scale and direction of the singular component.
Figure 7: Net change in multi-task loss when each singular component of a target task vector is individually added to a model merged with full-rank vectors from other tasks. (Left) Loss change induced by adding scaled component $\mathbf{s}_{i}=\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$. (Right) Loss change induced by adding component $\bar{\mathbf{s}}_{i}=\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$.
Figure 8: Change in multi-task loss caused by joint interaction of adding two singular components $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ together. We fix $\mathbf{s}_{i}$ as the top-1 component of task $i$ and sweep through the full indices of task $j$. (Left) Joint interaction measured with normalized components. (Right) Joint interaction measured with original scaled components.
Figure 9: (a) Loss change induced by adding component $\tilde{\mathbf{s}}_{i}=\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$. (b) $\Delta L_{\text{joint}}$ from adding $\tilde{\mathbf{s}}_{i}=\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ and $\tilde{\mathbf{s}}_{j}=\sigma_{j}\mathbf{u}_{j}\mathbf{v}_{j}^{\top}$ simultaneously. (c), (d) Values of the each terms in the Taylor expansion of the loss along singular directions (Equation 7): first-order term $\sigma_{i}\nabla L^{\top}\mathbf{s}_{i}$ (c) and second-order term $\sigma_{i}^{2}\mathbf{s}_{i}^{\top}H\mathbf{s}_{i}$ (d), respectively.
Singular Value Magnitude vs. Directional Alignment.
We now investigate the primary cause driving the interference in the top components. To decouple the effects of spectral scale and direction, we compare loss changes when adding (i) $\mathbf{s}_{i}=\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ (Figure 7 left) and (ii) $\bar{\mathbf{s}}_{i}=\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ (Figure 7 right), where singular values are removed. When we exclude the effect of the magnitude, the distribution of the interfering components becomes nearly uniform along the spectral indices, indicating that direction alone does not strongly distinguish the interference caused by top and bottom components. However, restoring singular values amplifies the change induced by top components due to their large $\sigma$: while beneficial ones now help more, misaligned ones cause larger increases in loss.
Alignment with Loss Curvature.
To further understand how singular directions interact with the loss landscape, we examine their alignment with the Hessian of the multi-task loss. Due to the heavy cost for calculating the full Hessian, we instead evaluate the quadratic term $\sigma_{i}^{2}\,\bar{\mathbf{s}}_{i}^{\top}\mathbf{H}\,\bar{\mathbf{s}}_{i}$ from Equation 7 using Hessian-vector products (HVP). Although this does not give explicit principal components of $\mathbf{H}$, it captures the aggregate alignment of each singular direction with high-curvature regions of the loss. Figure 9 (d) plots this quadratic term. Interestingly, we observe a distinct pattern compared to the total loss change: top singular components tend to exhibit positive curvature ($\bar{\mathbf{s}}_{i}^{\top}\mathbf{H}\,\bar{\mathbf{s}}_{i}>0$), while many bottom components show negative values. This suggests that the Hessian does introduce a directional effect—top components are more aligned with upward curvature than bottom components.
However, this curvature effect is not the primary driver of interference in the top range. In Figure 9 (c), we plot the first-order term $\bar{\sigma}\,\nabla L^{\top}\bar{\mathbf{s}}_{i}$. The first-order term exhibits behavior consistent with the actual loss change in Figure 9 (a): interfering directions ($\nabla L^{\top}\bar{\mathbf{s}}_{i}>0$) appear across the spectrum rather than being concentrated at the top indices. Taken together, these results indicate that while the Hessian alignment introduces a mild effect—top components tend to lie in regions of upward curvature—the first-order term dominates the overall loss change, and its interfering directions are distributed broadly across indices. Therefore, when comparing spectral scale and directional effects, the magnitude of the singular values remains the more dominant factor in shaping the top-range interference pattern observed in Section 3.
Joint Interaction Effects.
Finally, we consider the change emerging from the joint addition of singular components. Consider simultaneously adding two singular components $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$, each from task $i$ and $j$, to $\theta$. Using a second-order Taylor expansion, the multi-task loss change $\Delta L=L(\theta+\mathbf{s}_{i}+\mathbf{s}_{j})-L(\theta)$ can be expressed as:
$\Delta L\approx\underbrace{\sum_{p\in\{i,j\}}\left(\sigma_{p}\nabla L^{\top}\mathbf{\bar{s}}_{p}+\frac{1}{2}\sigma_{p}^{2}\mathbf{\bar{s}}_{p}^{\top}\mathbf{H}\mathbf{\bar{s}}_{p}\right)}_{\text{Individual Terms}}+\underbrace{\sigma_{i}\sigma_{j}(\mathbf{\bar{s}}_{i}^{\top}\mathbf{H}\mathbf{\bar{s}}_{j})}_{\text{Joint Term}}.$ (8)
The first two terms correspond to the loss changes from adding each component individually, as in Equation 7, and are independent of the other component. The joint interaction arises from the last term, which is a Hessian inner product between two directions. As in the single-component case, this interaction is governed by both the directional alignment of two singular vectors with respect to the loss curvature and the size of the singular values. Empirically, we quantify this joint interaction by measuring:
$\Delta L_{\text{joint}}=L(\theta+\mathbf{s}_{i}+\mathbf{s}_{j})-(L(\theta+\mathbf{s}_{i})+L(\theta+\mathbf{s}_{j})-L(\theta)),$ (9)
which isolates the additional loss change that occurs only when $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ are added together.
We fix $\mathbf{s}_{i}$ as the top-1 component from task $i$, and sweep over all singular components of task $j$ to measure this interaction. We observe a similar trend as in the single-component analysis (Figure 8); the directional interaction does not show a strong dependence on the index of $\mathbf{s}_{j}$, but interactions involving top components become much larger after accounting for their singular value. The joint effect is most amplified when both $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ originate from the top range, where their large singular values quadratically scale the interaction term. Therefore, joint interaction reinforces the same conclusion as the single-component addition: the risk of interference arises from misaligned directions being amplified by large singular values, especially for those prominent in the top range.
Summary.
Overall, we decomposed the loss change into first-order ($\nabla L^{\top}\bar{\mathbf{s}}_{i}$), quadratic ($\bar{\mathbf{s}}_{i}^{\top}\mathbf{H}\,\bar{\mathbf{s}}_{i}$), and joint ($\bar{\mathbf{s}}_{i}^{\top}\mathbf{H}\,\bar{\mathbf{s}}_{j}$) terms. We found that only the quadratic term shows a mild concentration of interfering components in the top range, while the first-order and joint terms produce interfering directions broadly across the spectrum. This suggests that the Hessian does introduce a directional effect, but this effect is secondary to the dominant impact of singular value scaling. In other words, fixed top-$k$ selection is suboptimal primarily because large singular values amplify any misaligned direction in the range, and this amplification effect is the main driver of top-range interference. This analysis also explains why AdaRank selects directions beyond top-$k$ (Figure 4 (a)): since bottom components often exhibit lower or negative curvature, they can provide useful updates with reduced interference. Consistent with this, Table 6 confirms that AdaRank’s primary gains come from pruning misaligned top components, while selecting non-top components yields smaller additional improvements.